Evaluate the given integral by changing to polar coordinates. 9(x + y) dA where R is the region that lies to the left of theCan you please solve both of these problems?
Answer #1

coordnete Il qix1 da alh 27424 2+4=4 - 4²=4 S=2 be the given integsed by chenging to palees heute fl 9 (x1) aA where R is theerlelhecete pe ves-syz da, Radny 1 202 2.42 €25,426 ere we went to evedute given inteysal by poleer for that we telke aarcoso